OpenStudy (anonymous):

Solve the following equation, answer as a reduced, mixed number. 7(x -2) + 3(x + 2) = 5(x - 6) x=___?

OpenStudy (anonymous):

7x-14+3x+6=5x-30 10x-8=5x-30 5x=22 x=22/5